OpenStudy (anonymous):

four cards are drawn from a pack of 52 cards.. what is the chance that.. (A)no two cards are of equal value.. (B)each belongs to a different suit..

OpenStudy (lgbasallote):

A) there are 13 different values so it will be 13/52 x 13/52 x 13/52 x13/52 got it?

OpenStudy (lgbasallote):

B) there are 4 different suits so it will be 4/52 x 4/52 x 4/52 x 4/52

OpenStudy (anonymous):

hmmm

OpenStudy (lgbasallote):

okay satellite is now gonna say im wrong :(

OpenStudy (anonymous):

the cards are not replaced

OpenStudy (lgbasallote):

oh...that makes a difference then

OpenStudy (lgbasallote):

A) it's gonna be decreasing so 13/52 x 12/51 x 11/50 x 10/49

OpenStudy (lgbasallote):

is that right @satellite73 ? sounds wrong...

OpenStudy (anonymous):

wrong ...:( try better @lgbasallote

OpenStudy (lgbasallote):

i think all the numerators should be 13

OpenStudy (anonymous):

each belong to a different suit \(\dbinom{52}{4}\) ways total, that is your denominator

OpenStudy (anonymous):

one from each suit, so \(\dbinom{13}{1}^4=13^4\) is your numerator

OpenStudy (lgbasallote):

ahh so all numerators are 13..

OpenStudy (anonymous):

no two card of equal value pick a card it is "different" with probability 1 then pick a second card. then the probability it is different value from the previous one is \(\frac{48}{51}\) pick a third probability it is different from first two is \(\frac{44}{51}\) etc

OpenStudy (anonymous):

???@satellite73 explain me in detail....plz..

OpenStudy (anonymous):

first one or second one?

OpenStudy (shubhamsrg):

a) no two cards are of equal value 52 card for the first choice, 52 - 4 since there are four the same value/number of cards. etc =52/52 *48/52 *44/52 *40/52 b) each belongs to a different suit. 52 card for the first choice ,since there are four suits 52/4=13. =52/52 * 39/52 * 26/52 * 13/52

OpenStudy (shubhamsrg):

i maybe wrong,please correct me if thats the case..

OpenStudy (anonymous):

i got it... @shubhamsrg u have done...a very slight mistake... 52/52 *48/51*44/50*40/49........since 1 card will be reduced each time... similarly...52/52 *39/51 *26/50*13/49...:)

OpenStudy (shubhamsrg):

1 card will be reduced when you take 1 by 1 here,,we have assumed that you take 4 at the sane time..